OpenStudy (anonymous):

Let ø: G--> G' be a homomorphism from the group G to the group G'. Prove that ø(a)= ø(b) if and only if ab^-1 exists in Ker(ø).

OpenStudy (anonymous):

The Kernel of G is the subset that is mapped to the identity of G' by ø. So I think you need to show that if ab^-1 is not in the Kernel, then ø(a) cannot = ø(b).

OpenStudy (anonymous):

yes, I know that I have to show two sides: the first side is : to show ø(a)= ø(b) and only if ab^-1 exists in Ker ( ø) Then I need to show that: If ab^-1 exists in Ker ( ø), then ø(a)= ø(b)

OpenStudy (anonymous):

If i am not mistaken though, ab^-1 IS in the kernel?

OpenStudy (anonymous):

only if ø(a)= ø(b) --- I have a hard time with proofs, because it seems obvious, right?

OpenStudy (anonymous):

if G is a homomorphism, then ø(ab) = ø(a)ø(b)

OpenStudy (anonymous):

...so ø(ab^-1) = ø(a)ø(b^-1) Isn't it true that if G is a homomorphism, (injective) that if ø(g) --> g' then there is no other g that is mapped to g'?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

...if ø(ab^-1) = ø(a)ø(b^-1) and ø(a)= ø(b), then ø(ab^-1) = ø(b)ø(b^-1) = ø(bb^-1) ...

OpenStudy (anonymous):

where did ø(a) go?

OpenStudy (anonymous):

= ø(b)

OpenStudy (anonymous):

okay so then is that equal to e'?

OpenStudy (anonymous):

Is the identity mapped to the identity?

OpenStudy (anonymous):

I was not given that information, butI thought that was given in a homomorphism

OpenStudy (anonymous):

However, I know that if ø is an isomorphism, then Ker ø={e}. and every isomorphism is a homomorphism